# Universal Transient Stability Analysis: A Large Language Model-Enabled Dynamics Prediction Framework

### v1

@@ 2.2 Transformer Attention Mechanism
The transformer architecture has demonstrated remarkable success in sequence modeling tasks due to its parallelizable computations and ability to capture long-range dependencies [vaswani2017attention]. At the core of this architecture is the self-attention mechanism. As illustrated in the bidirectional attention mechanism presented in Figure 1(a), the process begins with an input sequence $\mathbf{X}=[\mathbf{x}(1),\mathbf{x}(2),\ldots,\mathbf{x}(T)]\in\mathbb{R}^{T\times d}$, where $T$ is the sequence length and $d$ is the feature dimension, the bidirectional attention mechanism first transforms the input into three different representations: queries ($\mathbf{Q}$), keys ($\mathbf{K}$), and values ($\mathbf{V}$):
@@
where $\text{head}_{i}=\text{Attention}(\mathbf{X}\mathbf{W}^{Q}_{i},\mathbf{X}\mathbf{W}^{K}_{i},\mathbf{X}\mathbf{W}^{V}_{i})$. $\mathbf{W}^{Q}_{i}\in\mathbb{R}^{d\times d_{k}/h}$, $\mathbf{W}^{K}_{i}\in\mathbb{R}^{d\times d_{k}/h}$, $\mathbf{W}^{V}_{i}\in\mathbb{R}^{d\times d_{v}/h}$ are parameter matrices for the $i$-th attention head, $\mathbf{W}^{O}\in\mathbb{R}^{d_{v}\times d}$ is the output projection matrix, and $h$ is the number of attention heads. By operating on the entire sequence, this bidirectional attention mechanism is well-suited for tasks like sequence translation and classification [vaswani2017attention].
Figure 1: Comparison of attention mechanisms and iterative prediction in TSA-LLM.
@@ 3.1.1 Channel Independence
where each $x_{i}(t)\in\mathbb{R}$ corresponds to a distinct physical channel (e.g., $\delta_{i}$ or $\omega_{i}$). To overcome the fixed input dimensionality and the significant variance between stable and unstable dynamics, channel independence as illustrated in Figure 2(a) decouples the multivariate sequences $\mathbf{x}(t)$ into a set of univariate series $\{x_{i}(t)\}_{i=1}^{n_{x}}$, which are then processed and predicted independently. This design renders the model architecture agnostic to the number of input variables and enables robust handling of diverse dynamics across heterogeneous systems, thereby enhancing both cross-system scalability.
@@ 3.1.2 Segmentation and Sample-wise Normalization
Each univariate sequence $x_{i}\in\mathbb{R}^{T}$ is first partitioned into input-target samples via a sliding window of length $L_{\text{sam}}=L_{\text{seq}}+L_{\text{pred}}$ with a unit stride. Each sample contains an input sequence of length $L_{\text{seq}}$ and its corresponding prediction target of length $L_{\text{pred}}$. The input portion, $x_{ij,\text{input}}\in\mathbb{R}^{L_{\text{seq}}}$, is then subjected to sample-wise normalization to ensure numerical stability, particularly for high-magnitude unstable trajectories:
@@
where the mean $\mu_{ij}$ and standard deviation $\sigma_{ij}$ are computed from $x_{ij,\text{input}}$ rather than the entire training set. This sample-wise normalization, as presented in Figure 2(b) not only manages the extreme values of OOS SGs but also obviates the need for separate normalization statistics for stable/unstable SGs, a common practice in other methods [ye2024use]. Crucially, our approach eliminates the requirement for an online stability classification that may introduce both processing delays and the risk of error [tan2025bayesian], thus guaranteeing a more robust and streamlined framework suitable for real-time application.
@@ 3.1.3 Temporal Patch
The final data processing step is temporal patch, designed to enhance both feature representation and modeling efficiency. Each normalized sample $x_{ij}$ is partitioned into $P$ overlapping patches of length $L_{p}$ with a stride $S$, yielding a new, more compact representation ${x}^{P}_{ij}\in\mathbb{R}^{P\times L_{p}}$ as illustrated in Figure 2(c). The number of patches $P$ is given by:
@@
where $\lfloor\cdot\rfloor$ is the floor function. This patch-based tokenization enables the model to capture rich, local dynamic patterns (e.g., transient oscillations) that are more informative than isolated time points. Concurrently, by strategically reducing the input sequence length from $L_{\text{seq}}$ to $P$, it is instrumental in mitigating the quadratic complexity ($\mathcal{O}(N^{2})$) of the attention mechanism [vaswani2017attention], thereby ensuring the computational tractability of long-sequence modeling.
Figure 2: Model structure and pipeline of TSA-LLM.
@@ 3.2 Augmented GPT Architecture for TSA
Predicting long-horizon power system dynamics directly from a short observation window is an ill-posed problem, as the limited input resolution often leads to large prediction errors [shen2025physics]. A more principled solution is to adopt an iterative, autoregressive paradigm. In this approach, the model generates the future trajectory one step at a time, where each new prediction is conditioned on the sequence of all prior states (both observed and previously generated) [zhao2022structure]. This process is mathematically equivalent to the fundamental task of generative LLMs. It can be formally expressed by factorizing the conditional probability of the future trajectory, which enforces strict temporal causality:
@@
This inherent similarity is our core motivation for introducing an LLM architecture. We therefore adopt the open-source GPT architecture [radford2019language], a model renowned for its strong performance in generative tasks within natural language processing field. The model’s strength lies in its causal attention mechanism, which is fundamentally designed for the conditional, step-by-step sequence generation that accurate long-horizon forecasting requires. The subsequent sections describe the architecture of TSA-LLM, which is based on the GPT framework.
@@ 3.2.1 Input Embedding
The input embedding layer transforms the patch-structured samples ${x}^{P}_{ij}\in\mathbb{R}^{P\times L_{p}}$ into the initial hidden state $\mathbf{z}_{ij}^{0}$. This process is defined as:
@@ 3.2.2 Transformer Block Architecture
The backbone of TSA-LLM is composed of a stack of $L$ T-blocks adapted from the GPT model, as illustrated in Figure 2(e). Each block performs a transformation that refines an input sequence $\mathbf{z}^{l-1}$ into an output sequence $\mathbf{z}^{l}$:
@@
The MHCA layer containing $h$ parallel causal attention heads, endows the model with the ability to dynamically focus on relevant past information, as illustrated in Figure 1(b). Within each head, $\mathbf{z}_{ij}^{0}$ is projected into queries ($\mathbf{Q}$), keys ($\mathbf{K}$), and values ($\mathbf{V}$) using (2) to compute causal attention scores:
@@
where the lower-triangular mask, $\mathcal{M}_{\text{causal}}$ is formulated as follows:
@@
Pivotal to this operation is the causal mask, $\mathcal{M}_{\text{causal}}$, which nullifies the influence of future tokens by setting their corresponding attention scores to -$\infty$. This architectural constraint is what makes the model inherently suitable for iterative, generative tasks, as it forces the model to generate predictions based solely on available historical data, mimicking online dynamics prediction [liu2024timer]. The outputs of the $h$ heads are subsequently concatenated and linearly transformed by $\mathbf{W}^{O}$ to yield the multi-head representation as shown in (13).
@@
In (15), $\mathbf{W}_{1}\in\mathbb{R}^{d\times d_{\mathrm{ff}}}$, $\mathbf{W}_{{2}}\in\mathbb{R}^{d_{\mathrm{ff}}\times d}$ are learnable parameters and $d_{\mathrm{ff}}$ is the hidden dimension. $\operatorname{GELU(\cdot)}$ activation provides a smooth and non-linear transformation, while the final projection $\textbf{W}_{2}$ maps features back to the input dimensionality $d$, facilitating the composition of deep transformer architectures through block stacking.
@@ 3.2.3 Output Projection
The output layer transforms the final latent representation from the transformer, $\mathbf{z}_{ij}^{L}\in\mathbb{R}^{P\times d}$, into the prediction sequence, as illustrated in Figure 2(f). The process begins by flattening the patch-level representations in a single vector ${z}_{ij}^{\text{flat}}=\operatorname{Flatten}({z}_{ij}^{L})\in\mathbb{R}^{Pd}$ to consolidate all temporal features. This flattened vector is then mapped to the desired prediction horizon, via a learnable linear projection:
${x}_{ij}^{\text{pred}}={z}_{ij}^{\text{flat}}\cdot\mathbf{W}^{\text{out}}+\mathbf{b}^{\text{out}}\in\mathbb{R}^{L_{\text{pred}}},$ (16)
where $\mathbf{W}^{\text{out}}\in\mathbb{R}^{Pd\times L_{\text{pred}}}$ and $\mathbf{b}^{\text{out}}\in\mathbb{R}^{L_{\text{pred}}}$ are the learnable weights and biases. To restore the prediction to its original physical scale, an inverse normalization operation is applied:
$\hat{x}_{ij,\text{target}}=\sigma_{ij}\cdot{x}_{ij}^{\text{pred}}+\mu_{ij},$ (17)
where the mean ($\mu_{ij}$) and standard deviation ($\sigma_{ij}$) are the same values stored from the input sample’s normalization step (see (6)). The resulting vector, $\hat{x}_{ij,\text{target}}\in\mathbb{R}^{L_{\text{pred}}}$, represents the final physical prediction for the target sequence of length.
@@ 3.3.1 Freeze-and-Finetune Strategy
To balance adaptation to the TSA task with the preservation of knowledge from pre-trained T-blocks, we propose a parameter-efficient fine-tuning strategy. In this approach, the core T-blocks are kept frozen, while only the parameters of task-specific components, namely the input embedding layer, the layer normalization, and the output projection head, are updated during fine-tuning as illustrated in Figure 2. This strategy is designed to preserve the robust, generic feature extraction capabilities that the T-blocks acquire during pre-training. As detailed in Appendix A, the T-blocks are inherently optimized to find self- and cross-alignments within input sequences during pre-training stage, making it a powerful, modality-agnostic sequence feature extractor. Freezing the T-blocks leverages this proven capability. The efficacy of this strategy is validated in our case study. The high feature stability and co-direction (detailed in Section 4.8) confirm that the generic features extracted by the pre-trained T-blocks are highly effective for power system dynamics sequences. The following sections detail the proposed two-stage fine-tuning approach, where TeaF trains on sequence segments with ground truth sequences as input, and SchS addresses trajectory-level learning with probabilistically mixed inputs of model predictions and ground truth.
@@
12:  $\hat{x}_{i}\leftarrow\text{IterativePredict}(\mathbf{F},x_{i})$
@@ 3.3.2 Teacher Forcing Stage
During the first TeaF stage, the proposed TSA-LLM, represented by $\operatorname{\mathbf{F}}(\cdot)$, is conditioned on the ground-truth input segments, $x_{ij,\text{input}}$, to generate a prediction for the target sequence (denoted as $\hat{{x}}_{ij,\text{target}}=\operatorname{\mathbf{F}}({x}_{ij,\text{input}})$). The model parameters are optimized by minimizing the mean squared error (MSE) loss, $\mathcal{L}_{\text{TeaF}}$, between the predicted output and the ground-truth target:
$\mathcal{L}_{\text{TeaF}}=\frac{1}{L_{\text{pred}}}\sum_{t=1}^{L_{\text{pred}}}\left(\hat{x}_{ij,\text{target}}(t)-x_{ij,\text{target}}(t)\right)^{2}.$ (18)
This TeaF stage, as presented in Algorithm 1, provides a robust model initialization that accelerates convergence in subsequent SchS phases while enabling the model to achieve high fidelity on less complex transient trajectories. It is noteworthy that for numerous existing approaches, their entire training process consists solely of this TeaF paradigm [zhao2022structure, ye2024use, liu2024timer]. However, the iterative prediction process employed in online TSA application, where the TSA-LLM recursively conditions on its own outputs as illustrated in Figure 1(c), presents a training-inference mismatch. The model, trained via TeaF on ground-truth data, suffers from this mismatch and is not equipped to handle its own prediction errors, leading to cumulative error propagation in complex unstable trajectories [lamb2016professor].

### v2

@@ 2.2 Transformer Attention Mechanism
The transformer architecture has become the dominant backbone of modern large language models [1, 18, 26, 27] and has demonstrated remarkable success in sequence modeling, owing to its parallelizable computation and ability to capture long-range dependencies [36, 10]. At the core of this architecture is the self-attention mechanism. As illustrated in the bidirectional attention mechanism presented in Fig. 1(a), the process begins with an input sequence $\mathbf{X}=[\mathbf{x}(1),\mathbf{x}(2),\ldots,\mathbf{x}(T)]\in\mathbb{R}^{T\times d}$, where $T$ is the sequence length and $d$ is the feature dimension, the bidirectional attention mechanism first transforms the input into three different representations: queries ($\mathbf{Q}$), keys ($\mathbf{K}$), and values ($\mathbf{V}$):
@@
where $\text{head}_{i}=\text{Attention}(\mathbf{X}\mathbf{W}^{Q}_{i},\mathbf{X}\mathbf{W}^{K}_{i},\mathbf{X}\mathbf{W}^{V}_{i})$. $\mathbf{W}^{Q}_{i}\in\mathbb{R}^{d\times d_{k}/h}$, $\mathbf{W}^{K}_{i}\in\mathbb{R}^{d\times d_{k}/h}$, $\mathbf{W}^{V}_{i}\in\mathbb{R}^{d\times d_{v}/h}$ are parameter matrices for the $i$-th attention head, $\mathbf{W}^{O}\in\mathbb{R}^{d_{v}\times d}$ is the output projection matrix, and $h$ is the number of attention heads. By operating on the entire sequence, this bidirectional attention mechanism is well-suited for tasks like sequence translation and classification [36, 10].
Fig. 1: Comparison of attention mechanisms and iterative pblackiction in TSA-LLM.
@@ 3.1.1 Channel Independence
where each $x_{i}(t)\in\mathbb{R}$ corresponds to a distinct state variable channel. Channel independence is formulated at the level of $\mathbf{x}(t)$: rather than treating $\mathbf{x}(t)$ as a fixed-dimensional vector sequence, we decompose it into a set of scalar sequences $\{x_{i}(t)\}_{i=1}^{n_{x}}$. Each channel is processed by an identical, shablack pblackictor after applying the same segmentation and temporal-patch tokenization. This channel-wise shablack modeling yields a dimension-agnostic input/output interface, and thus eliminates the need to reparameterize the model to match the system-specific value of $n_{x}$.
For clarity, we use the rotor angle as a concrete example and denote $\delta(t)=[\delta_{1}(t),\ldots,\delta_{n_{g}}(t)]^{\top}$, where $n_{g}$ is the number of SGs (e.g., $n_{g}=10$ for the IEEE 39-bus system and $n_{g}=54$ for the IEEE 118-bus system). Across heterogeneous systems, $n_{g}$ varies with the network configuration, and so does the dimension of $\delta(t)$. Channel independence addresses this mismatch by mapping $\delta(t)\in\mathbb{R}^{n_{g}}$ to $\{\delta_{i}(t)\}_{i=1}^{n_{g}}$ and applying the same shablack pblackictor to each channel. Consequently, TSA-LLM does not rely on any fixed system-level input/output dimension and can be directly reused across systems with different generator numbers without architecture blackefinition.
Moreover, TSA-LLM keeps the model interface trajectory-centric: the inputs are measublack/simulated post-fault trajectories without concatenating explicit topology/parameter tables such as $Y$-bus, line parameters, or generator parameters. In this setting, topology- and parameter-dependent coupling effects are learned from the response patterns embedded in the transient trajectories, which preserves a unified interface that is convenient to reuse across heterogeneous systems.
@@ 3.1.2 Segmentation and Sample-wise Normalization
Each univariate sequence $x_{i}\in\mathbb{R}^{T}$ is first partitioned into input-target samples via a sliding window of length $L_{\text{sam}}=L_{\text{seq}}+L_{\text{pblack}}$ with a unit stride. Each sample contains an input sequence of length $L_{\text{seq}}$ and its corresponding pblackiction target of length $L_{\text{pblack}}$. The input portion, $x_{ij,\text{input}}\in\mathbb{R}^{L_{\text{seq}}}$, is then subjected to sample-wise normalization to ensure numerical stability, particularly for high-magnitude unstable trajectories:
@@
where the mean $\mu_{ij}$ and standard deviation $\sigma_{ij}$ are computed from $x_{ij,\text{input}}$ rather than the entire training set. This sample-wise normalization, as presented in Fig. 2(b), not only manages the extreme values of OOS SGs but also obviates the need for separate normalization statistics for stable/unstable SGs, a common practice in other methods [40]. Crucially, our approach eliminates the requirement for an online stability classification that may introduce both processing delays and the risk of error [34], thus guaranteeing a more robust and streamlined framework suitable for real-time application.
@@ 3.1.3 Temporal Patch
The final data processing step is temporal patch, designed to enhance both feature representation and modeling efficiency. Each normalized sample $x_{ij}$ is partitioned into $P$ overlapping patches of length $L_{p}$ with a stride $S$, yielding a new, more compact representation ${x}^{P}_{ij}\in\mathbb{R}^{P\times L_{p}}$ as illustrated in Fig. 2(c). The number of patches $P$ is given by:
@@
where $\lfloor\cdot\rfloor$ is the floor function. This patch-based tokenization enables the model to capture rich, local dynamic patterns (e.g., transient oscillations) that are more informative than isolated time points. Concurrently, by strategically blackucing the input sequence length from $L_{\text{seq}}$ to $P$, it is instrumental in mitigating the quadratic complexity ($\mathcal{O}(N^{2})$) of the attention mechanism [36], thereby ensuring the computational tractability of long-sequence modeling.
Fig. 2: Model structure and pipeline of TSA-LLM.
@@ 3.2 Augmented GPT Architecture for TSA
Pblackicting long-horizon post-fault dynamics from a short observation window is intrinsically ill-posed, because multiple future trajectories may be consistent with the same limited history and small modeling errors can rapidly accumulate over long rollouts [29]. To align with the time-evolution nature of post-fault transients, we adopt an iterative autoregressive formulation: the model generates the future trajectory step-by-step, and each pblackiction is conditioned on the observed history and all previously generated states [43]. This yields a strictly causal factorization of the conditional distribution, which is mathematically consistent with the next-token generation principle in generative LLMs:
@@
Motivated by this alignment, we adopt a Generative Pre-trained Transformer (GPT) backbone [23] and adapt it for transient trajectory generation. In particular, GPT’s causal self-attention naturally enforces temporal causality and provides a strong sequence prior for robust long-horizon autoregressive rollout. The subsequent sections describe the resulting GPT-based architecture of TSA-LLM.
@@ 3.2.1 Input Embedding
The input embedding layer transforms the patch-structublack samples ${x}^{P}_{ij}\in\mathbb{R}^{P\times L_{p}}$ into the initial hidden state $\mathbf{z}_{ij}^{0}$. This process is defined as:
@@ 3.2.2 Transformer Block Architecture
The backbone of TSA-LLM is composed of a stack of $L$ T-blocks adapted from the GPT model, as illustrated in Fig. 2(e). Each block performs a transformation that refines an input sequence $\mathbf{z}^{l-1}$ into an output sequence $\mathbf{z}^{l}$:
@@
The MHCA layer containing $h$ parallel causal attention heads, endows the model with the ability to dynamically focus on relevant past information, as illustrated in Fig. 1(b). Within each head, $\mathbf{z}_{ij}^{0}$ is projected into queries ($\mathbf{Q}$), keys ($\mathbf{K}$), and values ($\mathbf{V}$) using (2) to compute causal attention scores:
@@
where the lower-triangular mask, $\mathcal{M}_{\text{causal}}$, is formulated as follows. Here, $r$ and $c$ denote the row and column indices (i.e., the query-token position and the key-token position) in the attention-score matrix.
@@
Pivotal to this operation is the causal mask, $\mathcal{M}_{\text{causal}}$, which nullifies the influence of future tokens by setting their corresponding attention scores to -$\infty$. This architectural constraint is what makes the model inherently suitable for iterative, generative tasks, as it forces the model to generate pblackictions based solely on available historical data, mimicking online dynamics pblackiction [19]. The outputs of the $h$ heads are subsequently concatenated and linearly transformed by $\mathbf{W}^{O}$ to yield the multi-head representation as shown in (13).
@@
In (15), $\mathbf{W}_{1}\in\mathbb{R}^{d\times d_{\mathrm{ff}}}$, $\mathbf{W}_{{2}}\in\mathbb{R}^{d_{\mathrm{ff}}\times d}$ are learnable parameters and $d_{\mathrm{ff}}$ is the hidden dimension. $\operatorname{GELU(\cdot)}$ activation provides a smooth and nonlinear transformation, while the final projection $\textbf{W}_{2}$ maps features back to the input dimensionality $d$, facilitating the composition of deep transformer architectures through block stacking.
@@ 3.2.3 Output Projection
The output layer transforms the final latent representation from the transformer, $\mathbf{z}_{ij}^{L}\in\mathbb{R}^{P\times d}$, into the pblackiction sequence, as illustrated in Fig. 2(f). The process begins by flattening the patch-level representations in a single vector ${z}_{ij}^{\text{flat}}=\operatorname{Flatten}({z}_{ij}^{L})\in\mathbb{R}^{Pd}$ to consolidate all temporal features. This flattened vector is then mapped to the desiblack pblackiction horizon, via a learnable linear projection:
${x}_{ij}^{\text{pblack}}={z}_{ij}^{\text{flat}}\cdot\mathbf{W}^{\text{out}}+\mathbf{b}^{\text{out}}\in\mathbb{R}^{L_{\text{pblack}}},$ (16)
where $\mathbf{W}^{\text{out}}\in\mathbb{R}^{Pd\times L_{\text{pblack}}}$ and $\mathbf{b}^{\text{out}}\in\mathbb{R}^{L_{\text{pblack}}}$ are the learnable weights and biases. To restore the pblackiction to its original physical scale, an inverse normalization operation is applied:
$\hat{x}_{ij,\text{target}}=\sigma_{ij}\cdot{x}_{ij}^{\text{pblack}}+\mu_{ij},$ (17)
where the mean ($\mu_{ij}$) and standard deviation ($\sigma_{ij}$) are the same values stoblack from the input sample’s normalization step (see (6)). The resulting vector, $\hat{x}_{ij,\text{target}}\in\mathbb{R}^{L_{\text{pblack}}}$, represents the final physical pblackiction for the target sequence of length $L_{\text{pblack}}$.
@@ 3.3.1 Freeze-and-Finetune Strategy
To balance adaptation to the TSA task with the preservation of knowledge from pre-trained T-blocks, we propose a parameter-efficient fine-tuning strategy. In this approach, the core T-blocks are kept frozen, while only the parameters of task-specific components, namely the input embedding layer, the layer normalization, and the output projection head, are updated during fine-tuning as illustrated in Fig. 2. This strategy is designed to preserve the robust, generic feature extraction capabilities that the T-blocks acquire during pre-training. As detailed in A, the T-blocks are inherently optimized to find self- and cross-alignments within input sequences during the pre-training stage, making them powerful, modality-agnostic sequence feature extractors. Freezing the T-blocks leverages this proven capability. The efficacy of this strategy is validated in our case study. The high feature stability and co-direction (detailed in Section 4.8) confirm that the generic features extracted by the pre-trained T-blocks are highly effective for power system dynamics sequences. The following sections detail the proposed two-stage fine-tuning approach, where TeaF trains on sequence segments with ground truth sequences as input, and SchS addresses trajectory-level learning with probabilistically mixed inputs of model pblackictions and ground truth.
@@
12:  $\hat{x}_{i}\leftarrow\text{IterativePblackict}(\mathbf{F},x_{i})$
@@ 3.3.2 Teacher Forcing Stage
During the first TeaF stage, the proposed TSA-LLM, represented by $\operatorname{\mathbf{F}}(\cdot)$, is conditioned on the ground-truth input segments, $x_{ij,\text{input}}$, to generate a pblackiction for the target sequence (denoted as $\hat{{x}}_{ij,\text{target}}=\operatorname{\mathbf{F}}({x}_{ij,\text{input}})$). The model parameters are optimized by minimizing the mean squablack error (MSE) loss, $\mathcal{L}_{\text{TeaF}}$, between the pblackicted output and the ground-truth target:
$\mathcal{L}_{\text{TeaF}}=\frac{1}{L_{\text{pblack}}}\sum_{t=1}^{L_{\text{pblack}}}\left(\hat{x}_{ij,\text{target}}(t)-x_{ij,\text{target}}(t)\right)^{2}.$ (18)
This TeaF stage, as presented in Algorithm 1, provides a robust model initialization that accelerates convergence in subsequent SchS phases while enabling the model to achieve high fidelity on less complex transient trajectories. It is noteworthy that for numerous existing approaches, their entire training process consists solely of this TeaF paradigm [43, 40, 19]. However, the iterative pblackiction process employed in online TSA application, where the TSA-LLM recursively conditions on its own outputs as illustrated in Fig. 1(c), presents a training-inference mismatch. The model, trained via TeaF on ground-truth data, suffers from this mismatch and is not equipped to handle its own pblackiction errors, leading to cumulative error propagation in complex unstable trajectories [16].